OpenStudy (anonymous):

Perform the indicated operation g(x) = x3 - 2 h(x) = 2x - 1

OpenStudy (anonymous):

h(x)=x2(2 means squared)-1 (a) h(2)= (2)^2-1 = 4-1 = 3 (b)h(-2)= (-2)^2-1 = 4-1 = 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u r welcome :)